# Cohomogeneity one $\mathsf{RCD}$-spaces

### v2

@@ 4.3. Topological Rigidity
Let $(X,d)$ be a geodesic space, and $G$ compact acting by isometries on $X$. Assume that the orbit space $(X^{\ast},d^{\ast})$ is to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$. Then:
  1. (a)
    When $X^{\ast}$ is homeomorphic to $[0,1]$, then $X$ is homeomorphic to the union of two cones bundles.
  2. (b)
    When $X^{\ast}$ is homeomorphic to $[0,\infty)$, then $X$ is homeomorphic to a cone bundle
  3. (c)
  4. (d)
    When $X^{\ast}$ is homeomorphic to $\mathbb{R}$, then $X$ is homeomorphic to ×$\mathbb{R}\times G/H$.
@@
We begin by proving (c), and (d). By Lemma 3.11 it follows that for any point $x\in X$, the isotropy $G_{x}$ equals a fix subgroup $H<G$. Moreover, by Theorem 2.26 we conclude that for any $x\in X$, the orbit through $x$ is equivariantly homeomorphic to $G/H$, i.e. $G(x)\cong G/H$. Then, as pointed out in the proof of [27, Corollary 4.9], any lift $\tilde{\gamma^{\ast}}\colon I\to X$ of any short enough geodesic $\gamma^{\ast}\colon I\to\mathbf{S}^{1}=X^{\ast}$ gives a trivalization. In the case (c), since any two points in $X$ have conjugated isotropy subgroups, we get the conclusion on the structure group of the bundle. The case (d) follows from the fact that $\mathbb{R}$ is contractible.
The cases (a) and (b) follow from Remark 3.1, Theorem 2.29, and the observation that the spaces $\pi^{-1}([-1,0])$ and $\pi^{-1}([0,1])$ are equivariantly homeomorphic to the cone bundles $G\times_{G_{-}}\operatorname{Cl}(S_{-})$ and $G\times_{G_{+}}\operatorname{Cl}(S_{+})$ respectively. Here $\pi\colon X\to X^{\ast}$ is the orbit quotient map, $G_{\pm}$ are the isotropy groups of the orbits $\pi^{-1}(\{\pm 1\})$, and $S_{\pm}$ are slices through points in $\pi^{-1}(\{\pm 1\})$. ∎
Remark 4.13.
Observe that Theorem 4.12 is valid when $G$ is a non-compact group acting properly, and such that the orbit space is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$.
With Theorem 4.12 we can compute the fundamental group of the geodesic spaces of cohomogeneity one, following [45].
Theorem 4.14.
Let $(X,d)$ be a geodesic space, and $G$ a compact Lie group acting by isometries, such that $X^{\ast}$ is isometric to $[-1,1]$. We also consider a tuple $(G,H,G_{+},G_{-})$ given by Theorem 3.14 and assume that the orbits $K_{\pm}/H$ are connected. Then the fundamental group of $X$ is isomorphic to
where $N_{\pm}=\ker\{(\rho_{\pm})_{\ast}\colon\pi_{1}(G/H)\to\pi_{1}(G/H)\}$ for the orbit projection maps $\rho_{\pm}\colon G/H\to G/G_{\pm}$ give by $\rho_{\pm}(gH):=gK_{\pm}$.
Proof.
We recall that the group tuple $(G,H,G_{+},G_{-})$ is associated to a geodesic $\gamma\colon[-1,1]\to X$ realizing the distance between the orbits. Set $x_{\pm}:=\gamma(\pm 1)$ and $x_{0}=\gamma(0)$. Then $x_{\pm}\in\pi^{-1}(\pm 1)$ and $x_{0}\in\pi^{-1}(0)$.
@@
We point out that the slices $S_{x_{\pm}}$ are contractible, and thus the tubular neighborhoods $\mathrm{Tub}(\pi^{-1}(\pm 1)\cong G\times_{G_{\pm}}S_{x_{\pm}}$ onto the orbits $G(x_{\pm})$. Moreover, the point $x_{0}$ goes to the points $x_{\pm}$ under this deformation retract. Thus we have that $U_{\pm}\simeq\mathrm{Tub}(\pi^{-1}(\pm 1)\simeq\pi^{-1}(\pm 1)\cong G/K_{\pm}$, and $U_{-}\cap U_{+}\simeq\pi^{-1}(0)\cong G/H$. Under this homotopy equivalences we have that $x_{0}$ is mapped to the classes $K_{\pm}\in G/K_{\pm}$ and $H\in G/H$. Thus we have the following commutative diagram:
@@
From Theorem 4.14 we obtain the following sufficient conditions for a compact geodesic cohomogeneity one space to be simply connected.
@@
Let $(X,d)$ be a geodesic space, and $G$ a compact Lie group acting by isometries, such that $X^{\ast}$ is isometric to $[-1,1]$. We also consider a tuple $(G,H,G_{+},G_{-})$ given by Theorem 3.14 and assume that the orbits and $K_{\pm}/H$ are connected. If the inclusions
@@
we have that $\ker(\rho_{\pm})_{\ast}=\mathrm{img}(i_{\pm})_{\ast}$. Then the conclusion follows from Theorem 4.14. ∎
@@ 5. Construction of cohomogeneity one $\mathsf{RCD}$-spaces
we say that it is a cohomogeneity one group diagram if the homogeneous spaces $G_{\pm}/H$ admit $G_{\pm}$-invariant Riemannian metrics with positive Ricci curvature.
@@
Theorem B.
Let $G$ be a compact Lie group and $G_{+}$, $G_{-}$, $H$ Lie subgroups of $G$ such that $(G,H,G_{+},G_{-})$ is a cohomogeneity one group diagram. Then, there exists an $\mathsf{RCD}(K,N)$-space $(X,dist,\mathfrak{m})$ admitting a cohomogeneity one action of $G$ by measure preserving isometries, such that the associated group diagram is $(G,H,G_{+},G_{-})$. Moreover, we have $K\leq 0$, and $N=\max\{n_{-}+1+D,n_{+}+1+D,n_{0}+1\}$, where $n_{\pm}=\dim(G_{\pm}/H)$, $n_{0}=\dim(G/H)$ and $D=\dim(G)$.
@@
In this section we will work with certain warped products over compact Riemannian manifolds. The distances of the warped products that we will be considering will always be defined on $[0,\infty)\times X/\sim$, where $(t,x)\sim(s,y)$ if and only if $t=s=0$. The warping function $f$ that we consider only vanish at $t=0$.
@@
Let $R>0,$ and consider $\gamma\colon[0,1]\rightarrow[0,\infty)\times_{f}X$ an admissible curve such that $\gamma_{0},\gamma_{1}\in[0,R]\times_{f}X$ but that $\gamma([0,1])\not\subset[0,R]\times_{f}X$. The goal is to find a curve $\tilde{\gamma}$ contained in $[0,R]\times_{f}X$ with length less than or equal to that of $\gamma$. We have $\gamma=(\alpha,\beta)$, where $\alpha\in AC([0,1],[0,\infty])$. Recall that the absolute value of an absolutely continuous function is still an absolutely continuous function. We define the admissible curve $\tilde{\gamma}=(\tilde{\alpha},\beta)$ where $\tilde{\alpha}(t):=\max\{0,R-|\alpha(t)-R|\}.$ Observe that $f^{2}\circ\tilde{\alpha}(t)\leq f^{2}\circ\alpha(t)$ for all $t\in[0,1]$ and that $|\dot{\tilde{\alpha}}_{t}|\leq|\dot{\alpha}_{t}|$ on a set of full measure. Therefore
[…]
and so we have the admissible curve we were looking for in $[0,R]\times X$. Hence we have the result. ∎
@@
We consider now warped products using the function $f_{A,\epsilon}$ as warping function, with the objective of showing pointed measured Gromov-Hausdorff convergence of these warped spaces to the warped space with respect to $f_{A}$. In view of the convexity granted by Proposition 5.3, and because of the specific properties of the $f_{A,\epsilon}$, henceforward we restrict ourselves to $[0,2/A]\times_{f_{A,\epsilon}}X$ and $[0,2/A]\times_{f_{A}}X$. Let $[t,x],[s,y]\in[0,2/A]\times X/\sim$ and consider the set
@@
$\beta\in AC([0,1],X),\>\gamma_{0}=[t,x],\gamma_{1}=[s,y]\big{\}}$
@@
Similarly to what happens with $f_{\epsilon}$, $f$ we have the following result.
@@
The length functionals $\ell_{\epsilon}$ and $\ell$ of the warped products $[0,2/A]\times_{f_{\epsilon}}X$ and $[0,2/A]\times_{f}X$, respectively, satisfy the following for all curves $\gamma\in\bar{\mathcal{A}}$:
@@
[…]
@@
[…]
@@
In order to prove the pointed measured Gromov-Hasudorff convergence, we show that the induced distances converge pointwise; thus we need to check the convergence of the infima of the lengths $\ell_{\epsilon}$ on $\bar{\mathcal{A}}$ to the infimum of the length $\ell$ on $\bar{\mathcal{A}}$. To do so we make use of the notion of $\Gamma$-convergence. The interested reader may consult [9] for further details.
@@
Recall that the length functionals are lower semi-continuous. Now, from Lemma 5.5 we know that $\ell_{\epsilon}$ are non-decreasing with respect to $\epsilon$. Then using [9, Remark 2.12 ii)] we get that
@@
The next ingredient we need is the equicoercivity of the family $\{\ell_{\epsilon}\}$ (see [9, Definition 2.9]), which we now prove.
@@
Let $\lambda\geq 0$ and consider the set $K_{\lambda}:=\{\gamma\mid\ell_{1}(\gamma)\leq 2\lambda\}$. As we are dealing only with curves in compact spaces, [13, Theorem 2.5.14] tells us that $K_{\lambda}$ is compact. We will only work with $0<\epsilon\leq 1$. Take $\gamma=(\alpha,\beta)\in\{\gamma\mid\ell_{\epsilon}(\gamma)\leq\lambda\}$. It is immediate to see that
[…]
@@
[…]
@@
[…]
@@
Note that for all $\epsilon>0$ we have that $f\leq 2f_{\epsilon}$. This is useful because it implies that the $f$-length $\ell$ is bounded from above by the $2f_{\epsilon}$-length, that is, $\ell\leq\ell_{2f_{\epsilon}}$. This will in turn yield that $d\leq d_{2f_{\epsilon}}$. Now, if $\gamma$ is an admissible curve then it has small $f_{\epsilon}$-length if and only if its $2f_{\epsilon}$-length is also small. Then, the induced distances $d_{\epsilon},d$ are equivalent. All this allows us to put on $\bar{\mathcal{A}}$ the sup norm induced by $d$ without any problem.
@@
Let $[t,x],[s,y]\in[0,1]\times X/\sim$. Then from Proposition 5.7 we have that the functionals $\ell_{\epsilon}\colon\bar{\mathcal{A}}\rightarrow[0,\infty]$ $\Gamma$-converge to the functional $\ell\colon\bar{\mathcal{A}}\rightarrow[0,\infty]$. Now Proposition 5.8 tells us that for $0<\epsilon\leq 1$ we have equicoercivity, so then we can apply the Fundamental Theorem of $\Gamma$-convergence (see [9, Theorem 2.10]) and obtain that:
@@
or equivalently that $d_{\epsilon}([t,x],[s,y])\rightarrow d([t,x],[s,y])$ as $\epsilon\rightarrow 0$. Since the points in $[0,1]\times X/\sim$ were chosen arbitrarily we have pointwise convergence of the distance functions. Now, it is easy to convince oneself that the following properties hold:
@@
To conclude just observe that $d_{\epsilon}$, $d$ are continuous functions on a compact space so then by Dini’s Theorem we have that the convergence must be uniform. ∎
@@
hold for all $[t,x],[s,y]\in[0,1]\times X/\sim.$ Observe that this last condition tells us that $\mathrm{Id}$ is a $\delta$-Gromov-Hausdorff approximation for sufficiently small $\epsilon$.
@@
We now prove that when we require $X$ to have small diameter, then near the boundary the metric on $[0,2/A]\times_{f}X$ is the product metric.
@@
Let $(X,d_{X})$ be a compact geodesic metric space with $\operatorname{diam}(X)<1/A$. Then there exists $\delta>0$ such that the metric of the warped product $[0,2/A]\times_{f}X$ restricted to $[2/A-\delta,2/A]\times_{f}X$ is the product metric.
@@
First we prove that for $t$ sufficiently close to $2/A$ the geodesics joining points in the set $\{t\}\times X$ are always contained in $[1/A,2/A]\times X$. Since the warping function $f$ equals $1$ for all $t\in[1/A,2/A]$ this will yield that $\{t\}\times X$ is actually totally geodesic.
Proceeding by contradiction we assume that for all $\delta>0$ there exists $t\in[2/A-\delta,2/A]$ such that there are points $[t,x],[t,y]\in\{t\}\times X$ such that a minimal geodesic $\gamma$ joining them is not contained in $[1/A,2/A]\times X$. First observe that this means that $d([t,x],[t,y])<\operatorname{diam}(X)<1/A$ with the diameter in terms of the distance in $X$. Define the points
@@
We will now prove that geodesics joining points in $[2/A-\delta,2/A]\times X$ must lie there. Again, we proceed by contradiction: take two points $[r,x],[s,y]$ with $r<s$, and a minimal geodesic $\gamma$ with $\gamma_{0}=[r,x],\gamma_{1}=[s,y]$ such that it is not contained in $[2/A-\delta,2/A]\times X$. Now consider $t_{0}:=\max\{t\,|\,\gamma_{t}\in\{r\}\times X\}.$ Notice that then $\gamma([t_{0},1])\subset[2/A-\delta,2/A]\times X$. The restriction of $\gamma$ to the interval $[0,t_{0}]$ can be reparametrized to obtain a geodesic between $[r,x]$ and $\gamma_{t_{0}}.$ However, the previous argument implies then that $\gamma([0,t_{0}])\subset[2/A-\delta,2/A]\times X$, giving the contradiction.
Finally, as we have proved that geodesics between points in $[2/A-\delta,2/A]\times X$ must lie in there, it follows that the metric must be the product one. ∎
We are now tasked with looking at the structure of the Sobolev space $W^{1,2}$ of the warped product we have considered. We use the tools developed in [34] where Gigli and Han give conditions to ensure that the minimal weak upper gradient of a Sobolev function $g$ in the warped product $[0,\infty)\times_{f}X$ can be described in terms on the minimal weak upper gradients of the functions $g^{(t)}:=g(t,\cdot)$ and $g^{(x)}:=g(\cdot,x)$.
@@
Let $(X,d,\mathfrak{m})$ be an infinitessimally Hilbertian space, and consider $f\colon[0,\infty)\rightarrow[0,\infty)$ to be a warping function such that $\{f=0\}$ is discrete and $f$ decays at least near its zeroes (see equations 3.14, 3.15 in [34]). Then the warped product $[0,\infty)\times_{f}X$ is infinitessimally Hilbertian.
@@
Take $h,g\in W^{1,2}([0,1]\times_{f}X)$. Then by [34, Propositions 3.10 and 3.12], we have that the minimal weak upper gradient of $h$ is given by:
[…]
And similarly for $g$. [36, Theorem 4.3.3] states that $W^{1,2}([0,1]\times_{f}X)$ being a Hilbert space is equivalent to a pointwise parallelogram law. So we just need to check that this parallelogram law is satisfied from the hypothesis that both $[0,1]$ and $X$ satisfy such a parallelogram law:
@@
With this definition in hand we recall the following result of Ketterer (building upon the work of Alexander and Bishop), regarding curvature bounds and $\mathcal{F}K$-concavity.

### v3

@@ 4.2. Topological Rigidity
Let $(X,d,\mathfrak{m})$ be an essentially non-branching space, and $G$ a compact Lie group acting by isometries, such that $X^{\ast}$ is isometric to [1,1][-1,1]. We also consider a tuple $(G,H,K_{+},K_{-})$ given by ?THM? LABEL:T:_Principal_isotropy_is_subgroup_of_singular_isotropy and assume that the orbits $K_{\pm}/H$ are connected. Then the fundamental group of $X$ is isomorphic to
where $N_{\pm}=\ker\{(\rho_{\pm})_{\ast}\colon\pi_{1}(G/H)\to\pi_{1}(G/H)\}$ for the orbit projection maps $\rho_{\pm}\colon G/H\to G/K_{\pm}$ give by $\rho_{\pm}(gH):=gK_{\pm}$.
@@
We recall that the group tuple $(G,H,G_{+},G_{-})$ is associated to a horizontal geodesic $\gamma\colon[0,1]\to X$. Set $x_{-}=\gamma_{0},x_{+}=\gamma_{1}$, and $x_{0}=\gamma_{1/2}$. Then $x_{\pm}\in\pi^{-1}(\pm 1)$ and $x_{0}\in\pi^{-1}(0)$.
@@
We point out that the slices $S_{x_{\pm}}$ are contractible, and thus the tubular neighborhoods $\mathrm{Tub}(\pi^{-1}(\pm 1))\cong G\times_{G_{\pm}}S_{x_{\pm}}$ onto the orbits $G(x_{\pm})$. Moreover, the point $x_{0}$ goes to the points $x_{\pm}$ under this deformation retract. Thus we have that $U_{\pm}\simeq\mathrm{Tub}(\pi^{-1}(\pm 1))\simeq\pi^{-1}(\pm 1)\cong G/K_{\pm}$, and $U_{-}\cap U_{+}\simeq\pi^{-1}(0)\cong G/H$. Under this homotopy equivalences we have that $x_{0}$ is mapped to the classes $K_{\pm}\in G/K_{\pm}$ and $H\in G/H$. Thus we have the following commutative diagram:
@@
From ?THM? LABEL:T:_fundamental_group we obtain the following sufficient conditions for a compact essentially non-branching cohomogeneity one space to be simply connected.
@@
Let $(X,d,\mathfrak{m})$ be an essentially non-branching space, and $G$ a compact Lie group acting by isometries, such that $X^{\ast}$ is isometric to [1,1][-1,1]. We also consider a tuple $(G,H,G_{+},G_{-})$ given by ?THM? LABEL:T:_Principal_isotropy_is_subgroup_of_singular_isotropy and assume that the orbits and $K_{\pm}/H$ are connected. If the inclusions
@@
we have that $\ker(\rho_{\pm})_{\ast}=\mathrm{img}(i_{\pm})_{\ast}$. Then the conclusion follows from ?THM? LABEL:T:_fundamental_group. ∎
@@ 5. Construction of cohomogeneity one $\mathsf{RCD}$-spaces
we say that it is a cohomogeneity one group diagram if the homogeneous spaces $G_{\pm}/H$ admit $G_{\pm}$-invariant Riemannian metrics with positive Ricci curvature when $\dim(G_{\pm}/H)\geq 2$, or have finite diameter when $\dim(G_{\pm}/H)=1$.
@@
Theorem LABEL:thm.rcd_space_from_group_diagram.
Let $G$ be a compact Lie group and $_{+}, GG_{-}$, $H$ Lie subgroups of $G$ such that $(G,H,G_{+},G_{-})$ is a cohomogeneity one group diagram. Then, there exists an $\mathsf{RCD}(K,N)$-space $(X,d,\mathfrak{m})$ admitting a cohomogeneity one action of $G$ by measure preserving isometries, such that the associated group diagram is $(G,H,G_{+},G_{-})$. Moreover, we have $K\leq 0$, and $N=\max\{n_{-}+1+D,n_{+}+1+D,n_{0}+1\}$, where $n_{\pm}=\dim(G_{\pm}/H)$, $n_{0}=\dim(G/H)$ and $D=\dim(G)$.
@@
In this section we will work with certain warped products over compact Riemannian manifolds. The distances of the warped products that we will be considering will always be defined on $0,\infty)\times X/\sim$, where $(t,x)\sim(s,y)$ if and only if $t=s=0$. The warping function $f$ that we consider only vanish at $t=0$.
@@
Let $R>0$, and consider $\gamma\colon[0,1]\rightarrow[0,\infty)\times_{f}X$ an admissible curve such that $\gamma_{0},\gamma_{1}\in[0,R]\times_{f}X$ but that $\gamma([0,1])\not\subset[0,R]\times_{f}X$. The goal is to find a curve $\tilde{\gamma}$ contained in $0,R]\times_{f}X$ with length less than or equal to that of $\gamma$. We have $\gamma=(\alpha,\beta)$, where $\alpha\in AC([0,1],[0,\infty])$. Recall that the absolute value of an absolutely continuous function is still an absolutely continuous function. We define the admissible curve $\tilde{\gamma}=(\tilde{\alpha},\beta)$ where $\tilde{\alpha}(t):=\max\{0,R-|\alpha(t)-R|\}$. Observe that $f^{2}\circ\tilde{\alpha}(t)\leq f^{2}\circ\alpha(t)$ for all $t\in[0,1]$ and that $|\dot{\tilde{\alpha}}_{t}|\leq|\dot{\alpha}_{t}|$ on a set of full measure. Therefore
$\ell(\tilde{\gamma})=\int^{1}_{0}\sqrt{|\dot{\tilde{\alpha}}_{t}|^{2}+f^{2}\circ\tilde{\alpha}(t)|\dot{\beta}_{t}|^{2}}dt\leq\int^{1}_{0}\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}dt=\ell(\gamma)$
and so we have the admissible curve we were looking for in $0,R]\times X$. Hence we have the result. ∎
@@
We consider now warped products using the function $f_{A,\epsilon}$ as warping function, with the objective of showing pointed measured Gromov-Hausdorff convergence of these warped spaces to the warped space with respect to $f_{A}$. In view of the convexity granted by ?THM? LABEL:prop.convexconelike, and because of the specific properties of the $f_{A,\epsilon}$, henceforward we restrict ourselves to $0,2/A]\times_{f_{A,\epsilon}}X$ and $0,2/A]\times_{f_{A}}X$. Let $t,x],[s,y]\in[0,2/A]\times X/\sim$ and consider the set
@@
$\beta\in AC([0,1],X),\>\gamma_{0}=[t,x],\gamma_{1}=[s,y]\big\}.$
@@
Similarly to what happens with $_{\epsilon}, ff$ we have the following result.
@@
The length functionals $\ell_{\epsilon}$ and $\ell$ of the warped products $0,2/A]\times_{f_{\epsilon}}X$ and $0,2/A]\times_{f}X$, respectively, satisfy the following for all curves $\gamma\in\bar{\mathcal{A}}$:
@@
$\int^{1}_{0}\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}_{\epsilon}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}dt\leq\int^{1}_{0}\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}dt.$
@@
$\lim_{\epsilon\rightarrow 0}\int^{1}_{0}\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}_{\epsilon}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}dt=\int^{1}_{0}\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}dt.$
@@
In order to prove the pointed measured Gromov-Hausdorff convergence, we show that the induced distances converge pointwise; thus we need to check the convergence of the infima of the lengths $\ell_{\epsilon}$ on $\bar{\mathcal{A}}$ to the infimum of the length $\ell$ on $\bar{\mathcal{A}}$. To do so we make use of the notion of $\Gamma$-convergence. The interested reader may consult [10] for further details.
@@
Recall that the length functionals are lower semi-continuous. Now, from ?THM? LABEL:lemma.increasing.lengths we know that $\ell_{\epsilon}$ are non-decreasing with respect to $\epsilon$. Then using [10, Remark 2.12 ii)] we get that
@@
The next ingredient we need is the equicoercivity of the family $\{\ell_{\epsilon}\}$ (see [10, Definition 2.9]), which we now prove.
@@
Let $\lambda\geq 0$ and consider the set $K_{\lambda}:=\{\gamma\mid\ell_{1}(\gamma)\leq 2\lambda\}$. As we are dealing only with curves in compact spaces, [14, Theorem 2.5.14] tells us that $K_{\lambda}$ is compact. We will only work with $0<\epsilon\leq 1$. Take $\gamma=(\alpha,\beta)\in\{\gamma\mid\ell_{\epsilon}(\gamma)\leq\lambda\}$. It is immediate to see that
$\int^{1}_{0}|\dot{\alpha}_{t}|dt\leq\lambda,\quad\int^{1}_{0}f_{\epsilon}\circ\alpha(t)|\dot{\beta}_{t}|dt\leq\lambda.$
@@
$\int^{1}_{0}f_{1}\circ\alpha(t)|\dot{\beta}_{t}|dt\leq\int^{1}_{0}f_{\epsilon}\circ\alpha(t)|\dot{\beta}_{t}|dt\leq\lambda.$
@@
$\int^{1}_{0}\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}_{1}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}dt\leq\int^{1}_{0}|\dot{\alpha}_{t}|dt+\int^{1}_{0}f_{1}\circ\alpha(t)|\dot{\beta}_{t}|dt\leq 2\lambda.$
@@
Note that for all $\epsilon>0$ we have that $f\leq 2f_{\epsilon}$. This is useful because it implies that the $f$-length $\ell$ is bounded from above by the $2f_{\epsilon}$-length, that is, $\ell\leq\ell_{2f_{\epsilon}}$. This will in turn yield that $d\leq d_{2f_{\epsilon}}$. Now, if $\gamma$ is an admissible curve then it has small $f_{\epsilon}$-length if and only if its $2f_{\epsilon}$-length is also small. Then, the induced distances $_{\epsilon}, dd$ are equivalent. All this allows us to put on $\bar{\mathcal{A}}$ the sup norm induced by $d$ without any problem.
@@
Let $t,x],[s,y]\in[0,1]\times X/\sim$. Then from ?THM? LABEL:prop.Gammaconvergence.lengths we have that the functionals $\ell_{\epsilon}\colon\bar{\mathcal{A}}\rightarrow[0,\infty]$ $\Gamma$-converge to the functional $\ell\colon\bar{\mathcal{A}}\rightarrow[0,\infty]$. Now ?THM? LABEL:prop.equicoercivity.lengths tells us that for $0<\epsilon\leq 1$ we have equicoercivity, so then we can apply the Fundamental Theorem of $\Gamma$-convergence (see [10, Theorem 2.10]) and obtain that:
@@
or equivalently that $d_{\epsilon}([t,x],[s,y])\rightarrow d([t,x],[s,y])$ as $\epsilon\rightarrow 0$. Since the points in $0,1]\times X/\sim$ were chosen arbitrarily we have pointwise convergence of the distance functions. Now, it is easy to convince oneself that the following properties hold:
@@
To conclude just observe that $_{\epsilon}, dd$ are continuous functions on a compact space so then by Dini’s Theorem we have that the convergence must be uniform. ∎
@@
hold for all $t,x],[s,y]\in[0,1]\times X/\sim$. Observe that this last condition tells us that $\mathrm{Id}$ is a $\delta$-Gromov-Hausdorff approximation for sufficiently small $\epsilon$.
@@
We now prove that when we require $X$ to have small diameter, then near the boundary the metric on $0,2/A]\times_{f}X$ is the product metric.
@@
Let $(X,d_{X})$ be a compact geodesic metric space with $\operatorname{diam}(X)<1/A$. Then there exists $\delta>0$ such that the metric of the warped product $0,2/A]\times_{f}X$ restricted to $2/A-\delta,2/A]\times_{f}X$ is the product metric.
@@
First we prove that for $t$ sufficiently close to $2/A$ the geodesics joining points in the set $\{t\}\times X$ are always contained in $1/A,2/A]\times X$. Since the warping function $f$ equals $1$ for all $t\in[1/A,2/A]$ this will yield that $\{t\}\times X$ is actually totally geodesic.
Proceeding by contradiction we assume that for all $\delta>0$ there exists $t\in[2/A-\delta,2/A]$ such that there are points $t,x],[t,y]\in\{t\}\times X$ such that a minimal geodesic $\gamma$ joining them is not contained in $1/A,2/A]\times X$. First observe that this means that $d([t,x],[t,y])<\operatorname{diam}(X)<1/A$ with the diameter in terms of the distance in $X$. Define the points
@@
We will now prove that geodesics joining points in $2/A-\delta,2/A]\times X$ must lie there. Again, we proceed by contradiction: take two points $[r,x],[s,y]$ with $r<s$, and a minimal geodesic $\gamma$ with $\gamma_{0}=[r,x],\gamma_{1}=[s,y]$ such that it is not contained in $2/A-\delta,2/A]\times X$. Now consider $t_{0}:=\max\{t\,|\,\gamma_{t}\in\{r\}\times X\}$. Notice that then $\gamma([t_{0},1])\subset[2/A-\delta,2/A]\times X$. The restriction of $\gamma$ to the interval $0,t_{0}]$ can be reparametrized to obtain a geodesic between $[r,x]$ and $\gamma_{t_{0}}$. However, the previous argument implies then that $\gamma([0,t_{0}])\subset[2/A-\delta,2/A]\times X$, giving the contradiction.
Finally, as we have proved that geodesics between points in $2/A-\delta,2/A]\times X$ must lie in there, it follows that the metric must be the product one. ∎
We are now tasked with looking at the structure of the Sobolev space $W^{1,2}$ of the warped product we have considered. We use the tools developed in [32] where Gigli and Han give conditions to ensure that the minimal weak upper gradient of a Sobolev function $g$ in the warped product $0,\infty)\times_{f}X$ can be described in terms on the minimal weak upper gradients of the functions $g^{(t)}:=g(t,\cdot)$ and $g^{(x)}:=g(\cdot,x)$.
@@
Let $(X,d,\mathfrak{m})$ be an infinitessimally Hilbertian space, and consider $f\colon[0,\infty)\rightarrow[0,\infty)$ to be a warping function such that $\{f=0\}$ is discrete and $f$ decays at least linearly near its zeroes (see equations 3.14, 3.15 in [32]). Then the warped product $0,\infty)\times_{f}X$ is infinitessimally Hilbertian.
@@
Take $h,g\in W^{1,2}([0,1]\times_{f}X)$. Then by [32, Propositions 3.10 and 3.12], we have that the minimal weak upper gradient of $h$ is given by:
$|\nabla h|^{2}(t,x)=f^{-2}(t)|\nabla h^{(t)}|^{2}(x)+|\nabla h^{(x)}|^{2}(t),\quad f^{N}dt\otimes\mathfrak{m}-\text{a.e.}$
And similarly for $g$. [34, Theorem 4.3.3] states that $W^{1,2}([0,1]\times_{f}X)$ being a Hilbert space is equivalent to a pointwise parallelogram law. So we just need to check that this parallelogram law is satisfied from the hypothesis that both $[0,1]$ and $X$ satisfy such a parallelogram law:
@@
With this definition in hand we recall the following result of Ketterer (building upon the work of Alexander and Bishop [2]), regarding curvature bounds and $\mathcal{F}K$-concavity.